OpenStudy (anonymous):

Find the smallest positive integer value of n for which 225n is a multiple of 540

OpenStudy (anonymous):

Well if u simplify 225n/540 u get 12n/5 it's much easier to figure out then and u get n=12

OpenStudy (perl):

small typo that should say 5n/12

Parth (parthkohli):

Step 1: find the lowest common multiple of both the numbers.

OpenStudy (skullpatrol):

$$\large 225 = 3^2 \cdot 5^2$$ and $$\large 540 = 2^2 \cdot 3^3 \cdot 5 $$If 225n is a multiple of 540, then $$\large 3^2 \cdot 5^2 \cdot n = 2^2 \cdot 3^3 \cdot 5 \cdot k$$ where k and n are integers (which can be written as products of primes). What could n and k be to balance this equation?

OpenStudy (skullpatrol):

Obviously, k = 5 on the RHS. That leaves $$n = 2^2 \cdot 3 = 12$$